OpenStudy (anonymous):

What is the domain of the function? y = √3x+3 x ≤ – 1 x > 1 x ≥ –1

Parth (parthkohli):

If we get -1, we might get confused because of complex numbers. So, we have only one option i.e., x > 1

jimthompson5910 (jim_thompson5910):

You cannot take the square root of a negative number. So 3x+3 must be 0 or positive. which means that, 3x+3 >= 0 3x+3-3 >= 0-3 3x >= -3 3x/3 >= -3/3 x >= -1 So the answer is choice C